### v1

# On the extension of a class of Hermite bivariate interpolation problems

Abstract
We characterize the sets of solvability for Hermite bivariate interpolation problems when the sum of multiplicities is at most , with the degree of the polynomial space. This result extends an earlier theorem (2000) by one of the authors concerning the case . The latter theorem, in turn, can be regarded as a natural generalization of a classical theorem of Severi (1921).
Key words: Fundamental polynomial, algebraic curve, -correct set, -solvable set.
Mathematics Subject Classification (2010):
primary: 41A05, 41A63; secondary 14H50.
1 Introduction
In this paper, we characterize the sets of solvability for Hermite bivariate interpolation problems when the sum of multiplicities is at most , where denotes the degree of the polynomial space (see Theorem 2.1). This extends a previous result (Theorem 1.20) that addresses the case of . The latter theorem can, in turn, be regarded as a natural generalization of a classical theorem of Severi (Theorem 1.21).
One should mention that the correctness of Hermite problems was considered first in algebraic geometry by M. Nagata [6]. In approximation theory the investigation was initiated by G. G. Lorentz and R. A. Lorentz [4].
It should be noted also that a similar study in the case of Lagrange bivariate interpolation is more advanced (see [1]).
1.1 Lagrange bivariate interpolation
Denote the space of bivariate polynomials of total degree with real or complex coefficients:
Let us fix a set of distinct nodes in the plane:
The problem of finding a polynomial satisfying the conditions
for a given data is called Lagrange interpolation problem.
Definition 1.1.
A set of nodes is called -solvable if for any data there exists a polynomial (not necessarily unique) , satisfying the conditions (1.2).
Denote by the restriction of on the set
A polynomial is called an -fundamental polynomial for if
We denote this -fundamental polynomial by
Let us mention that sometimes we call -fundamental a polynomial satisfying the conditions
since it equals a non-zero constant times
Definition 1.2.
A set of nodes is called -independent if each node has an -fundamental polynomial. Otherwise, it is called -dependent.
If a set is -independent then the following Lagrange formula gives a solution to the problem (1.2):
Thus we obtain:
Proposition 1.3.
A node set is -independent if and only if it is -solvable.
Fundamental polynomials are linearly independent. Therefore a necessary condition of -independence, or -solvability, is
Definition 1.4.
A set of nodes is called -correct if for any data there exists a unique polynomial , satisfying the conditions (1.2).
A necessary condition of -correctness is:
For the sets of nodes satisfying the above condition we have
Proposition 1.5.
Assume that Then the following conditions are equivalent:
The set is -correct.
The set is -solvable.
A plane algebraic curve is the zero set of some bivariate polynomial of degree To simplify notation, we shall use the same letter, say , to denote the polynomial and the curve given by the equation . In particular, by we denote a linear polynomial from and the line defined by the equation Similarly by we denote a quadratic polynomial from and the conic defined by the equation
1.2 Hermite bivariate interpolation
Hermite interpolation problem is determined by a pair where is a set of nodes given by (1.1) and is a scheme, i.e., a sequence of multiplicities of the nodes associated with the degree
As a rule we assume that the multiplicities in the scheme are non-increasing:
Denote
In the problem the interpolation parameters at a given node are the value of a polynomial and its partial derivatives up to the order
The problem of finding a polynomial satisfying the above conditions, for a data
is called Hermite interpolation problem.
We express briefly the conditions (1.4) as follows
Nodes whose multiplicity is equal to one are called simple. At these nodes only the value of a polynomial is interpolated in the conditions (1.4).
It is sometimes convenient to allow nodes of multiplicity in the problem In fact, such a node can be ignored because there is no interpolation condition.
Definition 1.6.
A problem is called -solvable if for any data there exists a polynomial (not necessarily unique) , satisfying (1.4).
A polynomial is called an -fundamental polynomial for a parameter if it is a solution to the problem (1.4) where
The above fundamental polynomial we denote by
Definition 1.7.
A problem is called -independent if for each parameter in (1.4) there is an -fundamental polynomial.
The Lagrange formula for the above problem giving a solution to the problem (1.4), takes the form:
From here we get:
Proposition 1.8.
A problem is -independent if and only if it is -solvable.
Denote
We have that the number of interpolation conditions at a node in (1.4) equals to Hence the total number of the conditions equals to
It is easily seen that fundamental polynomials are linearly independent. Also we have that Therefore a necessary condition of -independence, or -solvability, of the problem is
A scheme satisfying the above condition we call -scheme.
A scheme satisfying the condition
we call -scheme.
Definition 1.9.
A problem is called -correct if for any data there exists a unique polynomial , satisfying the conditions (1.4).
A necessary condition of -correctness is:
A scheme satisfying the above condition we call an exact scheme.
Below we have the analog of Proposition 1.5 for the Hermite interpolation.
Proposition 1.10.
Assume that a scheme is exact. Then the following conditions are equivalent:
The problem is -correct.
The problem is -solvable.
Proposition 1.11.
Assume that in the problem the scheme is -scheme. Then there exists
Indeed, note that is the number of interpolation conditions in (1.4) and each condition can be considered as a linear equation with respect to unknown coefficients of Thus we get a linear homogeneous system, where the number of equations is less than the number of unknowns. As we know such systems have non-trivial solutions.
In the simplest case the above proposition implies
Lemma 1.12.
For any five points, there exists a conic that passes through all of them.
For a -scheme we define the exact scheme
Accordingly, for a set we define
where is a set of simple points.
Proposition 1.13 ([5], lemma 2.1).
The problem is -correct, for some if and only if the problem is -solvable.
For the sake of completeness we present:
Proof.
The direction “only if” is evident. For the direction “if”, in view of Proposition 1.10, it is enough to prove that there exists a (simple) node such that the problem where and is -solvable. For this, in view of -solvability of we need just an -fundamental polynomial of the simple node
It is easily seen that the polynomial is the desired -fundamental polynomial of ∎
Let and
Remark 1.14 ([3], rem.1.14).
If the problem is -solvable, then so is the problem with sufficiently close to
Definition 1.15.
Let be a problem, where is given by (1.1), and be a curve of degree Then we define the scheme as follows
where if and otherwise.
Denote by the number of nodes from on the curve counted with the multiplicities from
Note that in what follows, this notation, as well as the previous definition, mainly will be used in cases where the curve is a line or a conic.
We will need the following well-known proposition (see e.g. [2], lemma 4):
Proposition 1.16.
Let be a scheme with and be a set of nodes with where is a line. Then
Set
The following proposition we will use for lines and (irreducible) conics.
Proposition 1.17.
Let be a scheme with non-increasing sequence of multiplicities. If the problem is -solvable for a set of nodes then for any curve of degree we have that
Proof.
Let the problem be -solvable for a set of nodes Assume, by way of contradiction, that for a curve of degree we have that
Then, Proposition 1.13 implies that
for some where is a set of simple points with and
Moreover, by making use of Remark 1.14 we may assume that the set contains no point from the curve
Consider the following scheme
The above condition on the curve in view of (1.7), means that
Then let us check that is -scheme. Indeed, by using (1.7) and (1.9), we get
Therefore, in view of Proposition 1.11, there exists a polynomial such that
Now one readily obtains for the polynomial that
This, in view of Proposition 1.10, contradicts the -correctness of the problem ∎
By applying Proposition 1.17 for the cases we obtain the following two necessary conditions for -solvability:
Corollary 1.18.
Let be a scheme with non-increasing sequence of multiplicities. If the problem is -solvable for a set of nodes then:
For any line we have that
For any conic we have that
In the sequel the condition (i) will be referred to as the -condition.
Similarly, the condition (ii) will be referred to as the -condition.
Note, that by taking into account the above condition (i) for the lines passing through any two nodes, we obtain
Corollary 1.19.
Let be a scheme with non-increasing sequence of multiplicities. If the problem is solvable for a set of nodes then
Now we are in a position to present the result (slightly reformulated), which we will extend.
Theorem 1.20 ([2], th. 9).
Let be a scheme, satisfying the condition
Then the problem is -solvable if and only if there is no line passing through more than nodes of
This theorem can be regarded as a natural generalization of the following classical theorem of Severi (1921).
Theorem 1.21 (Severi [7], th. 9).
Let be a scheme. The problem is solvable for any node set if and only if
From Theorem 1.20 we get
Corollary 1.22.
For any scheme with non-increasing sequence of multiplicities, we have that
Indeed, consider the problem where in no three nodes are collinear. Then according to Theorem 1.20 the problem is -solvable. Hence is -scheme.
Proposition 1.23.
Let be a scheme with non-increasing sequence of multiplicities such that
Suppose is a set of nodes with where is an irreducible conic. Then we have that
Proof.
Consider the following scheme:
We have that
Note that
The above inequality, in view of irreducibility of implies that the -condition is satisfied. Thus, in view of Theorem (1.20), we obtain that the problem is -solvable. Similarly we obtain that the problem is -solvable.
Now consider the following linear spaces:
Note that In view of … we have that
Thus we obtain that ∎
Below we consider the case of above proposition. Note that the cases are not possible.
Remark 1.24.
Let be a scheme with non-increasing sequence of multiplicities such that
Then we readily get that implies that While implies that
Both these schemes are exact. According to the above proof the problems are also solvable, provided with any irreducible conic. The latter condition in this case means that no three nodes are collinear. Thus the problems are correct:
This holds also with the scheme replaced by Therefore Proposition 1.23 is valid also in the case since the statement reduces to
Remark 1.25.
Note that according to Corollary 1.22 for any scheme with non-increasing sequence of multiplicities, we have that the conditions imply that is -scheme.
In the main result of this paper below we consider instead the condition which, in combination with does not imply that is -scheme. Indeed, for a counterexample, we just add a fifth multiplicity to the schemes and of the above remark.
This is why in the next theorem, unlike Theorem 1.20, we have the condition: is -scheme.
2 The result
In this section we will prove the following
Theorem 2.1.
Let be a -scheme satisfying the condition
Then the problem is -solvable if and only if
there is no line passing through more than nodes of ,
there is no conic passing through more than nodes of
Proof.
The direction “only if” follows from Corollary 1.18. Let us prove the direction “if”.
Note that the above condition (ii) means that there is no conic passing through all the nodes of
In view of Theorem 1.20 we can assume, in what follows, that
From now on suppose also that is a scheme with non-increasing sequence of multiplicities:
In view of Corollary 1.19 we have that
We can assume also that
Indeed, the only - scheme, in the case is the Taylor scheme which does not satisfy (2.2). While if then the conditions (2.2) and (2.4) imply that which is not - scheme.
We will apply induction on the degree of the scheme
Step 1. It is easily seen that if then there is no -scheme satisfying the condition (2.2).
In the case there is only one - scheme satisfying the conditions (2.2) and (2.5). According to the condition (ii) of Theorem, the six simple nodes of the set do not belong to a conic. Since is exact scheme, by Proposition 1.5 (iii), we get that is -correct.
Now, assuming the statement holds for all degrees we will prove it for where
Step 2. Assume that there is a line which passes through nodes of i.e., the following condition is satisfied:
Consider the scheme (see (1.6))
From (1.7), in view of the condition (2.6), we have that
Now let us check that is - scheme. Indeed, by using (2.7), we get
The condition (2.1) for readily follows from the fact that passes through at least two nodes of
Then let us verify that the -condition is satisfied with respect to the problem i.e., there is no line containing nodes of Assume conversely that there is such a line Then evidently and, in view of the relation
we obtain that contains nodes from Since contains nodes both from and we conclude that the line does not intersect at a node of Therefore, the (reducible) conic passes through nodes of which contradicts the condition (ii) of Theorem.
Finally let us verify the -condition for the problem Assume, by way of contradiction, that there is a conic containing nodes of In view of (2.9) this means that passes through all the nodes with
Note that if passes through at least three nodes of then in (2.9) we would obtain and hence the -condition is satisfied. It remains to consider the case where passes through exactly two nodes, say, In this case, in view of (2.6), we have that According to (2.5) we get that Hence we have that Thus passes through all the nodes of which is contradiction.
Therefore, the problem is -solvable by virtue of the induction hypothesis.
Now, Proposition 1.13 implies that there is a set consisting of simple nodes, such that the problem
where
and the nodes of correspond to the last ones in
Here, by making use of Remark 1.14, we can also assume that the set contains no node belonging to the line
Now consider the problem where
Let us check first that is exact scheme. Indeed, in view of (2.7), we have that
Next, let us prove that the problem is -correct. For this, let us assume, in view of Proposition 1.10, (iii), that where and prove that
By the choice of the set and the condition (2.6) we have that
Now, Proposition 1.16 implies that where
Now, in view of (2.10) and Proposition 1.10, we get and hence
Thus, the problem is -correct, which implies evidently that is -solvable. ∎
From now on, in view of Step 2, we can assume that the -condition is satisfied for the scheme which implies that
Step 3. Assume that there is a line passing through at least nodes of Consider the following scheme (see (1.6))
We have that
On the other hand, in view of Step 2, the -condition holds. Thus according to Theorem 1.20 we obtain that the problem is -solvable.
Now, Proposition 1.13 implies that there is a set consisting of simple nodes, such that the problem
where
Again, by making use of Rremark 1.14, we can assume also that the set contains no node belonging to the line
Then denote by a set of simple nodes in
Consider the problem where and
Let us check first that is exact scheme. Indeed, we have that
Next, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.10, (iii), that where and prove that
In view of the choice of the sets and we have that
Therefore, Proposition 1.16 implies that where
Now, in view of (2.13) and Proposition 1.10, we get and hence
Thus, the problem is -correct, which implies that is -solvable. ∎
From now on, in view of Step 3, we may assume that there are no three collinear nodes in the set
Step 4. First note that the condition (ii) of Theorem, in view of Lemma 1.12, implies that
Consider a conic passing through the nodes Since no three nodes are collinear the conic is irreducible.
Note that may not satisfy the condition (2.3). Thus, to apply Theorem 1.20, let us verify that
Indeed, if then and, in view of (2.12), this is evident.
Now, assume conversely, that the inequality holds for some
Then, in view of (2.3), we have that and therefore
On the other hand we have that if Hence we have that which contradicts the condition (2.1).
Thus, according to Theorem 1.20, the problem is -solvable.
Now, Proposition 1.13 implies that there is a set consisting of simple nodes, such that the problem
where
By making use of Remark 1.14 we can assume that the set contains no node belonging to the conic
Next, denote by a set of simple nodes in
Consider the problem where and
Let us check first that is exact scheme. Indeed, we have that
Next, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.10, (iii), that where and prove that
By taking into account the choice of the sets and we have that Therefore, Proposition 1.23 implies that where
Now, in view of (2.15) and Proposition 1.10, we get and hence
Thus, the problem is -correct, which implies that is -solvable. ∎