### v2

# On the extension of a class of Hermite multivariate interpolation problems

Abstract
We characterize the sets of solvability for Hermite multivariate interpolation problems when the sum of multiplicities is at most , with the degree of the polynomial space. This result extends an earlier theorem (2000) by one of the authors concerning the case . The latter theorem, in turn, can be regarded as a natural extension of a classical Theorem of Severi (1921).
Key words: Hermite interpolation, fundamental polynomial, algebraic curve, -correct set, -solvable set.
Mathematics Subject Classification (2010):
primary: 41A05, 41A63; secondary 14H50.
1 Introduction
In this paper, we characterize the sets of solvability for Hermite multivariate interpolation problems when the sum of multiplicities is at most , where denotes the degree of the polynomial space (see Theorem 2.1). This result extends a previous one (Theorem 1.18) that addresses the case of . The latter theorem can, in turn, be regarded as a natural extension of a classical Theorem of Severi (Theorem 1.19).
One should mention that the correctness of Hermite problems was considered first in algebraic geometry by M. Nagata [7]. In approximation theory the investigation was initiated by G. G. Lorentz and R. A. Lorentz [4], [5].
It should be noted also that a similar study in the case of Lagrange bivariate interpolation is more advanced (see [1]).
1.1 Hermite multivariate interpolation
Hermite interpolation by multivariate algebraic polynomials is considered. The interpolation parameters are the values of a function and its partial derivatives up to some orders at given points.
Denote for and the multi-index
Set for convenience
The space of polynomials in variables of total degree is denoted by
The coordinates of points and coefficients of polynomials can be taken either real or complex, thus leading to a real or complex interpolation problem, respectively.
Consider a set of distinct interpolation points
Consider also a scheme, i.e., a sequence of multiplicities of the points associated with the degree of interpolating polynomials
As a rule we assume that the multiplicities in the scheme are non-increasing:
In Hermite interpolation problem where is the dimension, the interpolation parameters at a given point are the value of a polynomial and its partial derivatives up to the order
The problem of finding a polynomial satisfying the above conditions, for a data
is called Hermite interpolation problem
From now on we express briefly the conditions (1.4) as follows
Points whose multiplicity is equal to one are called simple. At these points only the value of a polynomial is interpolated in the conditions (1.4).
It is sometimes convenient, for example in proofs, to allow points of multiplicity in the problem In fact, such a point can be ignored because there is no interpolation condition.
Definition 1.1.
A problem is called -solvable if for any data there exists a polynomial (not necessarily unique) , satisfying (1.4).
A polynomial is called an -fundamental polynomial for a parameter if it is a solution to the problem (1.4) where
The above fundamental polynomial we denote by
Let us mention that sometimes we call -fundamental a polynomial, which is a solution to the problem (1.4) where since it equals a non-zero constant times
Definition 1.2.
A problem is called -independent if for each parameter in (1.4) there exists an -fundamental polynomial.
The Lagrange formula for the above problem giving a solution to the problem (1.4), takes the form:
From here we get:
Proposition 1.3.
A problem is -independent if and only if it is -solvable.
We have that the number of interpolation conditions at a point in (1.4) equals to (see 1.1). Hence the total number of the conditions equals to
It is easily seen that fundamental polynomials are linearly independent. Therefore a necessary condition of -independence, or -solvability, of the problem is
A scheme satisfying the above condition we call -scheme.
A scheme satisfying the condition
we call -scheme.
Definition 1.4.
A problem is called -correct if for any data there exists a unique polynomial , satisfying the conditions (1.4).
A necessary condition of -correctness is:
A scheme satisfying the above condition we call exact scheme.
Proposition 1.5.
Assume that a scheme is exact. Then the -correctness of the problem is equivalent to each of the following conditions:
The problem is -solvable.
Proof.
Indeed, each of the conditions in (1.4) can be considered as a linear equation with respect to unknown coefficients of Thus for exact scheme the conditions (1.4) reduce to a square linear system (see (1.6)). Now, the condition (i) means that the linear system is consistent for any right-hand side values, while the condition (ii) means that the homogeneous linear system has a single trivial solution. It is well known that in the case of a square system each of these properties implies that the system has a unique solution for any right-hand side values. The later property means that the problem is -correct. ∎
Proposition 1.6.
Assume that in the problem the scheme is -scheme. Then there exists
Proof.
Indeed, in this case the conditions (1.4) reduce to a linear homogeneous system, where the number of equations is less than the number of unknowns. As we know such systems have non-trivial solutions. ∎
Thus, for the -scheme we obtain
Corollary 1.7.
For any five points in the plane, there exists a conic that passes through all of them.
For a -scheme we define the associated exact scheme
Accordingly, for a set we define
where is a set of simple points.
Proposition 1.8 ([6], Lem.2.1).
The problem is -correct, for some if and only if the problem is -solvable.
For the sake of completeness we present:
Proof.
The direction “only if” is evident. For the direction “if”, in view of Proposition 1.5(i), it is enough to prove that there exists a (simple) point such that the problem where and is -solvable. For this, in view of the -solvability of we need just to have an -fundamental polynomial of the simple point
Then, it is easily seen that the polynomial is the desired -fundamental polynomial of ∎
Denote
Corollary 1.9.
Suppose that the problem is -solvable. Then we have that
Proof.
Remark 1.10 ([3], rem.1.14).
Let and
If the problem is -solvable, then so is the problem with sufficiently close to
If the problem is -solvable, with all belonging to an -dimensional linear subspace then so is the problem where the projection of on is
Until the end of this subsection, we study the bivariate case.
In this case we use the following simplified notation:
In this case instead of (1.1) it is more convenient to use the notation
Then the equality (1.5) and the second equality of (1.2) become:
A plane algebraic curve is the zero set of some bivariate polynomial of degree To simplify notation, we shall use the same letter, say , to denote the polynomial and the curve given by the equation . In particular, by we denote a linear polynomial from and the line defined by the equation Similarly, by we denote a quadratic polynomial from and the conic defined by the equation
Definition 1.11.
Let be a problem, where is a set given by (1.3), and be a plane curve of degree Then we define the scheme as follows
where if and otherwise.
Denote by the number of points from on the curve counted with the multiplicities from
Note that this notation, as well as the previous definition, mainly will be used in the cases where the curve either is a line or a conic.
For denote
We will need the following well-known proposition:
Proposition 1.12.
Let be a scheme with and be a set of points with where is a line. Then
where and
Corollary 1.13.
Suppose is a problem, where is -scheme. Suppose also is a line with Then the problem is -solvable if and only if the problem is -solvable, where
1.2 Some results in the multivariate case
Set
The following proposition will be used for lines and conics.
Proposition 1.14.
Let be a scheme. If the problem is -solvable for a set of points then for any plane curve of degree we have that
Proof.
First consider the case Let the problem be -solvable for a set of points Assume, by way of contradiction, that for a curve of degree we have that
Then, Proposition 1.8 implies that
for some where is a set of simple points with and
Moreover, by making use of Remark 1.10 we may assume that the set contains no point from the curve
Now, consider the following scheme
The above condition on the curve in view of (1.10), means that
Then let us check that is -scheme. Indeed, by using (1.10) and (1.12), we obtain
Therefore, in view of Proposition 1.6, there exists a polynomial such that
Now we readily get that the polynomial satisfies the condition:
Now consider the case of general dimension: Assume that does not satisfy the condition, i.e., there is a plane curve passing through more than of its points. Consider the respective subproblem restricted to the interpolation points belonging to the curve As we proved above, this subproblem is not -solvable in the case of Therefore the original problem also is not -solvable. ∎
By applying Proposition 1.14 for the cases we obtain the following two necessary conditions for -solvability:
Corollary 1.15.
Let be a scheme. If the problem is -solvable for a set of points then:
For any line we have that
For any conic we have that
In the sequel the condition (i) will be referred to as the -condition.
Similarly, the condition (ii) will be referred to as the -condition.
Definition 1.16.
A scheme is called regular if there exists a set of points such that the problem is -independent.
By taking into account the above condition (i) for the lines passing through any two points, and the condition (ii) for the conics passing through any five points, one readily obtains
Corollary 1.17.
Let be a scheme with non-increasing sequence of multiplicities. If is regular then we have that
Now we are in a position to present the result, which we are going to extend. Actually, it is slightly reformulated (see [2], 2nd par. after the proof of Thm.9).
Theorem 1.18 ([2], Thm.9).
Let be a scheme, satisfying the condition
Then the problem is -solvable if and only if there is no line passing through more than points of
This theorem can be regarded as a natural extension of the following classical Theorem of Severi (1921).
Theorem 1.19 (Severi [8], Thm.9).
Let be a scheme. Then the problem is -solvable for any point set if and only if
Until the end of the subsection, we study the bivariate case again.
From Theorem 1.18 we get
Corollary 1.20.
Suppose that a scheme with non-increasing sequence of multiplicities, satisfies the condition: Then is regular if only if
Proof.
For denote
Next proposition is an extension of Proposition 1.12:
Proposition 1.21.
Let be a scheme with non-increasing sequence of multiplicities such that
Suppose is a set of points with where is an irreducible conic. Then we have that
where and
Proof.
Consider the following scheme:
We have that
Note that
The above inequality, in view of irreducibility of implies that the -condition is satisfied for the problem .
Thus, in view of Theorem (1.18), we obtain that is -solvable. Similarly we obtain that the problem is -solvable.
Now, consider the following linear spaces:
Note that Since the problems and are solvable, we get from Corollary 1.9, that
Thus we obtain that ∎
Below we consider the case of above proposition. Note that the cases are not possible.
Remark 1.22.
Let be a scheme with non-increasing sequence of multiplicities such that
Then implies that While implies that
Now, let Note that is exact. The proof of Proposition 1.21 implies that the problem is solvable, provided that with any irreducible conic. The latter condition in this case means that no three points of are collinear. Thus the problem is correct:
Note that also (see the above proof) is correct, since if then Therefore Proposition 1.21 is valid also in the case and the statement (1.13), i.e., reduces to
Remark 1.23.
Note that, according to Corollary 1.20, for any scheme with non-increasing sequence of multiplicities, we have that the conditions imply that is -scheme.
In the main result of this paper below we consider instead the condition which, in combination with does not imply that is -scheme. Indeed, for a counterexample, we just add a fifth multiplicity to the scheme of the above remark.
This is why, in the following theorem we have an additional condition, compared to Theorem 1.18: is -scheme.
2 The main result
In this section we will prove the following
Theorem 2.1.
Let be a -scheme satisfying the condition
Then the problem is -solvable if and only if:
There is no line passing through more than points of
There is no conic passing through more than points of
Proof.
The direction “only if” follows from Corollary 1.15. Let us prove the direction “if”. First we consider the (main) case
Note that the above condition (ii) means that there is no conic passing through all the points of
In view of Theorem 1.18 we can assume, in what follows, that
From now on suppose also that is a scheme with non-increasing sequence of multiplicities:
By using the condition (i) for the line passing through we get
We can assume also that
Indeed, assume that Then the only -scheme is the the Taylor scheme But the latter does not satisfy (2.1). While if then the conditions (2.1) and (2.3) imply that which is not -scheme.
We will apply induction on the degree of the scheme
Step 1. It is easily seen that if then there is no -scheme satisfying the condition (2.1).
In the case there is only one -scheme satisfying the conditions (2.1) and (2.4). According to the condition (ii) of Theorem, the six simple points of the set do not belong to a conic. Since is exact scheme, we obtain from Proposition 1.5(ii), that is -correct.
Now, assuming the statement holds for all degrees we will prove it for where
Step 2. Assume that there is a line which passes through points of i.e., the following condition is satisfied:
Consider the scheme (see (1.9))
From (1.10), in view of the condition (2.5), we have that
Next, we want to apply the induction hypothesis for the problem For this let us check first that is -scheme. Indeed, by using (2.6), we get
The condition (2.1) for readily follows from the fact that passes through at least two points of
Then let us verify that the -condition is satisfied with respect to the problem i.e., there is no line containing points of Assume conversely that there is such a line Then evidently and, in view of the relation
we obtain that contains points from Since contains points both from and we conclude that the line does not intersect at a point of Therefore, the (reducible) conic passes through points of which contradicts the condition (ii) of Theorem.
Finally let us verify the -condition for the problem Assume, by way of contradiction, that there is a conic containing points of In view of (2.7) this means that passes through all the points with
Note that if passes through at least three points of then in (2.7) we would obtain and hence the -condition is satisfied. It remains to consider the case where passes through exactly two points of , say, and In this case, in view of (2.5), we have that According to (2.4) we get that Hence we have that Thus passes through all the points of which is contradiction.
Therefore, the problem is -solvable by virtue of the induction hypothesis.
Now, Proposition 1.8 implies that there is a set consisting of simple points (see (1.7) and (1.8)), such that the problem
where
Here, by making use of Remark 1.10, we can also assume that the set contains no point belonging to the line
Now consider the problem where
Let us check first that is exact scheme. Indeed, in view of (2.6), we have
Next, let us prove that the problem is -correct. For this, let us assume, in view of Proposition 1.5(ii), that where and prove that
By the choice of the set and the condition (2.5) we have that
Then, Proposition 1.12 implies that where
Now, in view of (2.8) and Proposition 1.5(ii), we get and hence
Thus, the problem is -correct, which implies evidently that is -solvable. ∎
From now on, in view of Step 2, we can assume that the -condition is satisfied for the scheme which implies that
Step 3. Assume that there is a line passing through at least points of Consider the following scheme (see (1.9))
We have that
On the other hand, in view of Step 2, the -condition holds. Thus according to Theorem 1.18 we obtain that the problem is -solvable.
Now, Proposition 1.8 implies that there is a set consisting of simple points, such that the problem
where
Again, by making use of Remark 1.10, we can assume also that the set contains no point belonging to the line
Then denote by a set of simple points in
Consider the problem where and
Let us check first that is exact scheme. Indeed, we have that
Next, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.5(ii), that where and prove that
In view of the choice of the sets and we have that
Therefore, Proposition 1.12 implies that where
Now, in view of (2.10) and Proposition 1.5(ii), we get and hence
Thus, the problem is -correct, which implies that is -solvable. ∎
From now on, in view of Step 3, we may assume that there are no three collinear points in the set
Step 4. First note that the condition (ii) of Theorem, in view of Corollary 1.7, implies that Then consider a conic passing through the points The conic is irreducible, since no three points are collinear.
Note that may not satisfy the condition (2.2). Thus, to apply Theorem 1.18, let us verify that
Indeed, if then and, in view of (2.9), this is evident.
Now, assume conversely, that the inequality holds for some Then, in view of (2.2), we have that and therefore
On the other hand we have that if Hence we have that which contradicts the condition (2.1).
Thus, according to Theorem 1.18, the problem is -solvable. Therefore, Proposition 1.8 implies that there is a set consisting of simple points, such that the problem
where
By making use of Remark 1.10 we can assume that the set contains no point belonging to the conic
Next, denote by a set of simple points in
Consider the problem where and
Let us check first that is exact scheme. Indeed, we have that
Now, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.5(ii), that where and prove that
By taking into account the choice of the sets and we have that Therefore, Proposition 1.21 implies that where
Then, in view of (2.11) and Proposition 1.5(ii), we get and hence
Thus, the problem is -correct, which implies that is -solvable. ∎
Step 5. It remains to consider the case of general dimension.
Assume that is a set of points in dimension satisfying the and -conditions with respect to the problem Then we can project it to a set of points in dimension, with the same property. Continuing in this way we will obtain a set of points in a plane, with the same property. Now, the problem is -solvable, since Theorem has already been proven in the bivariate case. From here, by using times Remark 1.10, we obtain that the original problem is -solvable too. ∎
A counterexample.
Suppose the scheme is the following Thus (2.12) takes place. Also we have that is -scheme:
Suppose the corresponding set of points is chosen as follows: The points all belong to an irredusible conic The points and belong to a line which does not contain any other point from
It is easily seen that the conditions (i) and (ii) of Theorem 2.1 are satisfied:
Let us verify that the Hermite problem is not -solvable.
Indeed, by applying Corollary 1.13 with respect to the line we get that the problem is -solvable if and only if the problem is -solvable, where and
Now, on the conic we have points of Hence, in view of Corollary 1.15(ii), the problem is not -solvable.
From Theorem 2.1 we get
Corollary 2.2.
Suppose that a -scheme with non-increasing sequence of multiplicities, satisfies the condition: Then is regular if and only if and
Proof.
Consider the following exact scheme This scheme satisfies the conditions of above corollary, with one exception: According to Corollary 1.17 the scheme is not regular.